# NyayaRAG: Realistic Legal Judgment Prediction with RAG under the Indian Common Law System

### v2

@@ Abstract
Legal Judgment Prediction (LJP) has emerged as a key area in AI for law, aiming to automate judicial outcome forecasting and enhance interpretability in legal reasoning. While previous approaches in the Indian context have relied on internal case content such as facts, issues, and reasoning, they often overlook a core element of common law systems, which is reliance on statutory provisions and judicial precedents. In this work, we propose NyayaRAG, a Retrieval-Augmented Generation (RAG) framework that simulates realistic courtroom scenarios by providing models with factual case descriptions, relevant legal statutes, and semantically retrieved prior cases. NyayaRAG evaluates the effectiveness of these combined inputs in predicting court decisions and generating legal explanations using a domain-specific pipeline tailored to the Indian legal system. We assess performance across various input configurations using both standard lexical and semantic metrics as well as LLM-based evaluators such as G-Eval. Our results show that augmenting factual inputs with structured legal knowledge significantly improves both predictive accuracy and explanation quality.
@@
Shubham Kumar Nigam1,5∗†  Balaramamahanthi Deepak Patnaik1∗  Shivam Mishra1∗ Ajay Varghese Thomas2∗Noel Shallum3Kripabandhu Ghosh4Arnab Bhattacharya1 1IIT Kanpur, India,  2SRM Institute of Science and Technology, India 3Symbiosis Law School Pune, India  4IISER Kolkata, India 5University of Birmingham, Dubai, United Arab Emirates {shubhamkumarnigam, bdeepakpatnaik2002, shivam1602m, ajaythomas.work, noelshallum} @gmail.com kripaghosh@iiserkol.ac.in,arnabb@cse.iitk.ac.in
@@ 1 Introduction
The application of artificial intelligence (AI) in legal judgment prediction (LJP) has the potential to transform legal systems by improving efficiency, transparency, and access to justice. This is particularly crucial for India, where millions of cases remain pending in courts, and decision-making is inherently dependent on factual narratives, statutory interpretation, and judicial precedent. India follows a common law system, where prior decisions (precedents) and statutory provisions play a central role in influencing legal outcomes. However, most existing AI-based LJP systems do not adequately replicate this fundamental feature of judicial reasoning.
@@
Previous studies such as malik-etal-2021-ildc; nigam-etal-2024-legal; nigam-etal-2025-nyayaanumana have focused on predicting legal outcomes using the current case document, including sections like facts, arguments, issues, reasoning, and decision. More recent efforts have narrowed the scope to factual inputs alone (nigam-etal-2024-rethinking; FactLegalLlama), yet these systems still operate in a vacuum, without considering how courts naturally rely on applicable laws and prior rulings. In reality, judges rarely decide in isolation; instead, they actively refer to relevant precedent and statutory law. To bridge this gap, we propose a framework that more closely mirrors actual courtroom conditions by explicitly incorporating external legal knowledge during inference.
@@
Moreover, in critical domains like finance, medicine, and law, decisions must be grounded in verifiable information. Experts in these domains cannot rely on opaque, black-box inferences, and they require systems that ensure factual consistency. Hallucinations, common in large generative models, can have severe consequences in legal decision-making. By retrieving and conditioning model responses on grounded sources such as applicable laws and precedent cases, Retrieval-Augmented Generation (RAG) offers a principled approach to mitigate hallucination and promote trustworthy outputs. Furthermore, RAG frameworks like ours can be flexibly integrated into existing legal systems without requiring the retraining of core models or the sharing of private or sensitive case data. This enhances user trust while allowing the legal community to benefit from AI without sacrificing transparency or data confidentiality.
@@
We introduce NyayaRAG, a Retrieval-Augmented Generation (RAG) framework for realistic legal judgment prediction and explanation in the Indian common law system. The term “NyayaRAG” is derived from two components: “Nyaya” meaning “justice” and “RAG” referring to “Retrieval-Augmented Generation”. Together, the name reflects our vision to build a justice-aware generation system that emulates the reasoning process followed by Indian courts, using facts, statutes, and precedents.
@@
Unlike prior models that operate purely on internal case content, NyayaRAG simulates real-world judicial decision-making by providing the model with: (i) the summarized factual background of the current case, (ii) relevant statutory provisions, (iii) top-k semantically retrieved previous similar judgments. This structure emulates how judges deliberate on new cases, consulting both textual statutes and prior judicial opinions. Through this design, we evaluate how Retrieval-Augmented Generation can help reduce hallucinations, promote faithfulness, and yield legally coherent predictions and explanations.
Our contributions are as follows:
@@
    Simulating Common Law Reasoning with LLMs: We show that LLMs guided by RAG and factual grounding can produce legally faithful explanations aligned with how real-world decisions are made under common law reasoning.
Our work moves beyond fact-only or self-contained models by replicating a more faithful legal reasoning pipeline aligned with Indian jurisprudence. We hope that NyayaRAG opens new directions for building interpretable, retrieval-aware AI systems in legal settings, particularly in resource-constrained yet precedent-driven judicial systems like India’s. For the sake of reproducibility, we have implemented our data set, code, and RAG-based pipeline through a GitHub repository [^1] https://github.com/ShubhamKumarNigam/NyayaRAG.
@@ 2 Related Work
Within the Indian legal ecosystem, the ILDC corpus (malik-etal-2021-ildc) laid the foundation for Court Judgment Prediction and Explanation (CJPE). This has been expanded in PredEx (nigam2024legaljudgmentreimaginedpredex), NyayaAnumana (nigam2024nyayaanumana), LegalSeg (nigam2025legalseg), TathyaNyaya (nigam2025tathyanyaya), and IBPS for bail prediction (srivastava2025ibps), which collectively focus on segmentation, factual reasoning, and judgment outcome explanation. Efforts such as nigam2024rethinking and 10.1145/3632754.3632765 emphasize fact-based LJP as a realistic alternative to full-text reasoning, aligning predictions more closely with early-phase decision-making. Other works explore semantic segmentation (malik2021semantic), legal QA (nigam2023legal), and structured document generation (nigam2025structured). These resources and models highlight the significance of domain-specific adaptation for Indian law, an underrepresented common law system.
Precedent integration is central to common law reasoning. Early precedent-aware models such as zhao-etal-2018-learning focused on Chinese law, while recent works like PLJP (wu2023precedent), LexKeyPlan (santosh2025lexkeyplan), and precedent-enhanced ECHR prediction (santosh2024incorporating) examine retrieval of historical judgments to guide predictions. These align with broader multilingual and cross-jurisdictional LJP studies (niklaus2021swiss; kapoor-etal-2022-hldc). While these efforts highlight the role of past cases, their settings differ significantly from India’s unique common law context where both statutory law and judicial precedents guide reasoning.
Retrieval-augmented pipelines are increasingly applied to legal NLP tasks. LegalBench-RAG (pipitone2024legalbench) and CLERC (hou2024clerc) introduced benchmarks for retrieval-augmented legal reasoning. Specialized systems like CBR-RAG (wiratunga2024cbr), Graph-RAG for norms (de2025graph), and LexKeyPlan (santosh2025lexkeyplan) explore structured retrieval, case-based reasoning, and keyphrase-driven planning. Applications range from legal assistants and dispute resolution (rafat2024ai; 10887211) to federated secure access architectures (amato2024optimizing). These works demonstrate the value of augmenting generative reasoning with statutory and precedent-based retrieval but primarily address civil law or international corpora.
Our work contributes to this line by contextualizing RAG for the Indian common law system, simulating how judges reason with case facts, statutory provisions, and precedents. While prior RAG frameworks in law (wiratunga2024cbr; hou2024clerc; wu2023precedent; santosh2024incorporating) establish the feasibility of retrieval-augmented pipelines, they have not been applied to Indian courts. By integrating factual, statutory, and precedential inputs in a unified framework, NyayaRAG extends the landscape of LJP toward realistic, explainable, and domain-grounded prediction and explanation.
@@ 3 Task Description
Figure 1 illustrates our Legal Judgment Prediction (LJP) pipeline enhanced with RAG. The pipeline begins with a full legal judgment document, which undergoes summarization to reduce its length and retain essential factual meaning. This is necessary because legal judgments tend to be long, and appending retrieved knowledge further increases the input size. Given limited model capacity and computational resources, we employ a summarization step (using Mixtral-8x7B-Instruct-v0.1) to create a condensed representation of both the input case and the retrieved legal context.
Prediction Task: Based on the summarized factual description $D$ and the retrieved top-$k$ (e.g., $k=3$) similar legal documents (statutes or precedents), the model predicts the likely court judgment. The prediction label $y\in\{0,1\}$ indicates whether the appeal is fully rejected (0) or fully/partially accepted (1). This binary framing captures the most common forms of judicial decisions in Indian appellate courts.
Explanation Task: Alongside the decision, the model is also required to generate an explanation that justifies its output. This explanation should logically incorporate the facts, cited statutes, and relevant precedents retrieved during the RAG process. This step emulates how judges provide reasoned opinions in written judgments.
By structuring the LJP task in this way, summarizing long documents and integrating retrieval-based augmentation, we study the effectiveness of RAG agents in producing judgments that are both faithful to legal reasoning and grounded in precedent and statute. The overall framework allows us to approximate a real-world decision-making environment within Indian courtrooms.
@@ 4 Dataset
Our dataset is designed to simulate realistic court decision-making in the Indian legal context, incorporating facts, statutes, and precedent, essential elements under the common law framework. This dataset enables exploration of Legal Judgment Prediction (LJP) in a Retrieval-Augmented Generation (RAG) setup.
@@ 4.1 Dataset Compilation
4.2 Single vs. Multi Partitioning
Another challenge in the SCI proceedings is the presence of cases containing multiple petitions under a single case identifier, where each petition may lead to distinct outcomes. To handle this, we partition the dataset into two subsets. The single partition represents binary classification tasks (accept/reject) and includes cases with a single petition or multiple petitions that share the same outcome. The multi partition, a superset of single, contains cases with multiple petitions leading to different decisions, introducing a more complex multi-label prediction setting (e.g., partially accepted outcomes). For these cases, if even one appeal within the group is accepted, the case is assigned the label accepted. Predicting such mixed outcomes remains challenging and will be further explored in future work.
@@ 4.3 Dataset Composition
The corpus supports multiple downstream pipelines, each focusing on specific judgment elements or legal context. Table 1 presents key statistics across different configurations, and an example breakdown is shown in the Appendix Table 7.
@@ 4.3.1 Case Text
Table 1: NyayaRAG Data Statistics.
@@
Table 2: Instruction prompt used with Mixtral-8x7B-Instruct-v0.1 for summarizing legal judgments.
@@ 4.3.4 Previous Similar Cases
This retrieval step enriches context with precedents that are semantically close, even if not cited, enhancing the legal realism of our setup.
@@ 4.3.5 Facts
Overall, our dataset is uniquely structured to test legal decision-making under realistic constraints, aligning with the Indian legal system’s reliance on factual narratives, statutory frameworks, and prior rulings.
@@ 5 Methodology
To simulate realistic judgment prediction and evaluate the role of RAG in enhancing legal decision-making, we design a modular experimental setup. This setup explores how different types of legal information, such as factual summaries, statutes, and precedents, affect model performance on the dual tasks of prediction and explanation. To ensure reproducibility and transparency, we detail the full experimental setup, including model configurations, training routines, and task-specific hyperparameters, in Appendix A. This includes separate subsections for the explanation generation (summarization) and legal judgment prediction tasks, outlining all relevant decoding strategies, optimization settings, and dataset splits used across our pipeline variants.
@@ 5.3 Inference Setup
We use the LLaMA-3.1 8B Instruct dubey2024llama model for all experiments in a few-shot prompting setup. Each input sequence, composed according to one of the pipeline templates, is paired with a relevant prompt. The model is required to output:
@@ 6 Evaluation Metrics
Pipeline Name Partition Accuracy Precision Recall F1-score
@@
Table 3: Performance of Various Pipelines on Binary and Multi-label Legal Judgment Prediction. The best result has been marked in bold.
@@
    Inter-Annotator Agreement (IAA): To further ensure the reliability and consistency of expert judgments, we computed multiple standard inter-rater agreement metrics. Specifically, Fleiss’ Kappa fleiss1971measuring was used to evaluate the overall agreement among multiple raters, Cohen’s Kappa cohen1960coefficient captured pairwise agreement while adjusting for chance, and the Intraclass Correlation Coefficient (ICC) shrout1979intraclass measured the reliability of continuous ratings across raters. In addition, Krippendorff’s Alpha krippendorff2018content provided a robust measure suitable for ordinal scales and missing data, while the Pearson Correlation Coefficient benesty2009pearson quantified the linear consistency of expert scores. The results demonstrated substantial agreement across these different measures, reinforcing the credibility of the evaluation and lending strong support to the robustness of our findings.
@@
Table 4: Comparison of Explanation Generation Across Different Legal Context Pipelines.

### v3

@@ Abstract
Legal Judgment Prediction (LJP) has emerged as a key area in AI for law, aiming to automate judicial outcome forecasting and enhance interpretability in legal reasoning. While previous approaches in the Indian context have relied on internal case content such as facts, issues, and reasoning, they often overlook a core element of common law systems – reliance on statutory provisions and judicial precedents. In this work, we propose NyayaRAG, a Retrieval-Augmented Generation (RAG) framework that simulates realistic courtroom scenarios by providing models with factual case descriptions, relevant legal statutes, and semantically retrieved prior cases. NyayaRAG evaluates the effectiveness of these combined inputs in predicting court decisions and generating legal explanations using a domain-specific pipeline tailored for the Indian legal system. We assess performance across various input configurations using both standard lexical and semantic metrics as well as LLM-based evaluators such as G-Eval. Our results show that augmenting factual inputs with structured legal knowledge significantly improves both predictive accuracy and explanation quality.
@@
Shubham Kumar Nigam1,5∗†  Balaramamahanthi Deepak Patnaik1∗  Shivam Mishra1∗ Ajay Varghese Thomas2∗Noel Shallum3Kripabandhu Ghosh4Arnab Bhattacharya1 1Indian Institute of Technology Kanpur, India  2SRM Institute of Science and Technology, India 3Symbiosis Law School Pune, India  4IISER Kolkata, India 5University of Birmingham, Dubai, United Arab Emirates {shubhamkumarnigam, bdeepakpatnaik2002, shivam1602m, ajaythomas.work, noelshallum}@gmail.com kripaghosh@iiserkol.ac.inarnabb@cse.iitk.ac.in
@@ 1 Introduction
The application of artificial intelligence (AI) in legal judgment prediction (LJP) has the potential to transform legal systems by improving efficiency, transparency, and access to justice. This is particularly crucial for India, where millions of cases remain pending in courts, and decision-making is inherently dependent on factual narratives, statutory interpretation, and judicial precedents. India follows a common law system, where prior decisions (precedents) and statutory provisions play a central role in influencing legal outcomes. However, most existing AI-based LJP systems do not adequately replicate this fundamental feature of judicial reasoning.
@@
Previous studies (malik-etal-2021-ildc; nigam-etal-2024-legal; nigam-etal-2025-nyayaanumana) have focused on predicting legal outcomes using the current case document, including sections like facts, arguments, issues, reasoning, and decision. More recent efforts have narrowed the scope to factual inputs alone (nigam-etal-2024-rethinking; FactLegalLlama); yet these systems still operate in a setting that is not real, and does not consider how courts naturally rely on applicable laws and prior rulings. In reality, judges rarely decide in isolation; instead, they actively refer to relevant precedents and statutory laws. To bridge this gap, we propose a framework that more closely mirrors actual courtroom conditions by explicitly incorporating external legal knowledge during inference.
@@
In critical domains like law, medicine and finance, decisions must be grounded in verifiable information. Experts in these domains cannot rely on opaque, black-box inferences and require systems that ensure factual consistency. Hallucinations, common in large generative models, can have severe consequences in legal decision-making. By retrieving and conditioning model responses on grounded sources such as applicable laws and precedent cases, the Retrieval-Augmented Generation (RAG) paradigm offers a principled approach to mitigate hallucination and promote trustworthy outputs. Furthermore, RAG frameworks like ours can be flexibly integrated into existing legal systems without requiring the retraining of core models or the sharing of private or sensitive case data. This enhances user trust while allowing the legal community to benefit from AI without sacrificing transparency or data confidentiality.
@@
Motivated by the above, we introduce NyayaRAG, a Retrieval-Augmented Generation (RAG) framework for realistic legal judgment prediction and explanation in the Indian common law system. The term “NyayaRAG” is derived from two components: “Nyaya” meaning “justice” and “RAG” referring to “Retrieval-Augmented Generation”. Together, the name reflects our vision to build a justice-aware generation system that emulates the reasoning process followed by Indian courts, using facts, statutes, and precedents.
@@
Unlike prior models that operate purely on internal case content, NyayaRAG simulates real-world judicial decision-making by providing the model with: (i) the summarized factual background of the current case, (ii) relevant statutory provisions, and (iii) retrieved top-k semantically similar previous judgments. This structure emulates how judges deliberate on new cases, consulting both textual statutes and prior judicial opinions. Through this design, we evaluate how RAG can help reduce hallucinations, promote faithfulness, and yield legally coherent predictions and explanations.
Our contributions in this paper are as follows:
@@
    Simulating Common Law Reasoning with LLMs: We show that LLMs guided by RAG and factual grounding can produce legally faithful explanations that are aligned with how real-world decisions are made under common law systems.
Our work moves beyond fact-only or self-contained models by replicating a more faithful legal reasoning pipeline aligned with Indian jurisprudence. We hope that NyayaRAG opens new directions for building interpretable, retrieval-aware AI systems in legal settings, particularly in resource-constrained yet precedent-driven judicial systems like India’s. We have shared our dataset, code, and RAG-based pipeline through a GitHub repository [^1] https://github.com/ShubhamKumarNigam/NyayaRAG.
@@ 2 Related Work
Within the Indian legal ecosystem, the ILDC corpus (malik-etal-2021-ildc) laid the foundation for Court Judgment Prediction and Explanation (CJPE). This has been expanded in PredEx (nigam2024legaljudgmentreimaginedpredex), NyayaAnumana (nigam2024nyayaanumana), LegalSeg (nigam2025legalseg), TathyaNyaya (nigam2025tathyanyaya), and IBPS for bail prediction (srivastava2025ibps), which collectively focus on segmentation, factual reasoning, and judgment outcome explanation. Efforts such as (nigam2024rethinking; 10.1145/3632754.3632765) emphasize fact-based LJP as a realistic alternative to full-text reasoning, aligning predictions more closely with early-phase decision-making. Other works explore semantic segmentation (malik2021semantic), legal QA (nigam2023legal), and structured document generation (nigam2025structured). These resources and models highlight the significance of domain-specific adaptation for Indian law, an underrepresented common law system.
Precedent integration is central to common law reasoning. Early precedent-aware models such as (zhao-etal-2018-learning) focused on Chinese law, while recent works like PLJP (wu2023precedent), LexKeyPlan (santosh2025lexkeyplan), and precedent-enhanced ECHR prediction (santosh2024incorporating) examine retrieval of historical judgments to guide predictions. These align with broader multilingual and cross-jurisdictional LJP studies (niklaus2021swiss; kapoor-etal-2022-hldc). While these efforts highlight the role of past cases, their settings differ significantly from India’s unique common law context where both statutory law and judicial precedents guide reasoning.
Retrieval-augmented pipelines are increasingly applied to legal NLP. LegalBench-RAG (pipitone2024legalbench) and CLERC (hou2024clerc) introduced benchmarks for retrieval-augmented legal reasoning. Specialized systems like CBR-RAG (wiratunga2024cbr), Graph-RAG for norms (de2025graph), and LexKeyPlan (santosh2025lexkeyplan) explore structured retrieval, case-based reasoning, and keyphrase-driven planning. Applications range from legal assistants and dispute resolution (rafat2024ai; 10887211) to federated secure access architectures (amato2024optimizing). These works demonstrate the value of augmenting generative reasoning with statutory and precedent-based retrieval but primarily address civil law or international corpora.
Our work contributes to this line by contextualizing RAG for the Indian common law system, simulating how judges reason with case facts, statutory provisions, and precedents. While prior RAG frameworks in law (wiratunga2024cbr; hou2024clerc; wu2023precedent; santosh2024incorporating) establish the feasibility of retrieval-augmented pipelines, they have not been applied to Indian courts. By integrating factual, statutory, and precedential inputs in a unified framework, NyayaRAG extends the landscape of LJP toward realistic, explainable, and domain-grounded legal tasks.
@@ 3 Task Description
Figure 1 illustrates our Legal Judgment Prediction (LJP) pipeline enhanced with RAG. The process begins with a full legal judgment document, which undergoes summarization to reduce its length and retain essential factual meaning. This is necessary because legal judgments tend to be long, and appending retrieved knowledge further increases the input size. Given limited model capacity and computational resources, we employ a summarization step (using Mixtral-8x7B-Instruct-v0.1) to create a condensed representation of both the input case and the retrieved legal context.
Prediction Task:
Based on the summarized factual description and the retrieved top-$k$ similar legal documents (statutes or precedents), the model predicts the likely court judgment. The prediction label $y\in\{0,1\}$ indicates whether the appeal is fully rejected (0) or fully/partially accepted ($1$). This binary framing captures the most common forms of judicial decisions in Indian appellate courts.
Explanation Task:
Alongside the decision, the model is also required to generate an explanation that justifies its output. This explanation should logically incorporate the facts, cited statutes, and relevant precedents retrieved during the RAG process. This step emulates how judges provide reasoned opinions in written judgments.
By structuring the LJP task in this manner, and by summarizing long documents and integrating retrieval-based augmentation, we study the effectiveness of RAG in producing judgments that are both faithful to legal reasoning and grounded in precedent and statute. The overall framework allows us to approximate real-world decision-making environments within Indian courtrooms.
@@ 4 Dataset
Our dataset is designed to simulate realistic court decision-making in the Indian legal context, incorporating facts, statutes, and precedent, which are essential elements under the common law framework. This dataset enables exploration of Legal Judgment Prediction (LJP) in a Retrieval-Augmented Generation (RAG) setup.
@@ 4.1 Dataset Compilation
4.2 Single versus Multi Partitioning
Another challenge in the SCI proceedings is the presence of cases containing multiple petitions under a single case identifier, where each petition may lead to distinct outcomes. To handle this, we partition the dataset into two subsets. The single partition represents binary classification tasks (accept/reject) and includes cases with a single petition or multiple petitions that share the same outcome. The multi partition contains cases with multiple petitions leading to different decisions, introducing a more complex multi-label prediction setting (e.g., partially accepted outcomes). For these cases, if even one appeal within the group is accepted, the case is assigned the label accepted. Predicting mixed outcomes remains challenging and will be further explored in future work.
@@ 4.3 Dataset Composition
The corpus supports multiple downstream pipelines, each focusing on specific judgment elements or legal context. Table 1 presents key statistics across different configurations. An example case is shown in Table 7 in Appendix.
@@ 4.3.1 Case Text
Table 1: NyayaRAG data statistics
@@
Table 2: Instruction prompt used with Mixtral-8x7B-Instruct-v0.1 for summarizing legal judgments
@@ 4.3.4 Previous Similar Cases
This retrieval step enriches context with precedents that are semantically close; even if not cited, they enhance the legal realism of our setup.
@@ 4.3.5 Facts
Overall, our dataset is uniquely structured to test legal decision-making under realistic constraints, aligning with the Indian legal system’s reliance on factual narratives, statutory frameworks, precedents, and prior rulings.
@@ 5 Methodology
To simulate realistic judgment prediction and evaluate the role of RAG in enhancing legal decision-making, we design a modular experimental setup. This setup explores how different types of legal information, such as factual summaries, statutes, and precedents affect model performance on the dual tasks of prediction and explanation. To ensure reproducibility and transparency, we detail the full experimental setup, including model configurations, training routines, and task-specific hyperparameters. The details are mentioned in Appendix A. This includes separate sections for explanation generation (summarization) and legal judgment prediction tasks, outlining all relevant decoding strategies, optimization settings, and dataset splits used across our pipeline variants.
@@ 5.3 Inference Setup
We use the LLaMA-3.1 8B Instruct dubey2024llama model for all experiments in a few-shot prompt setup. Each input sequence, composed according to one of the pipeline templates, is paired with a relevant prompt that asks to output:
@@ 6 Evaluation Metrics
Pipeline Partition Accuracy Precision Recall F1-score
@@
Table 3: Performance of various pipelines on Binary and Multi-label LJP task (best results in bold)
@@
    Inter-Annotator Agreement (IAA): To further ensure the reliability and consistency of expert judgments, we computed multiple standard inter-rater agreement metrics. Specifically, Fleiss’ Kappa fleiss1971measuring was used to evaluate the overall agreement among multiple raters, Cohen’s Kappa cohen1960coefficient captured pairwise agreement while adjusting for chance, and the Intraclass Correlation Coefficient (ICC) shrout1979intraclass measured the reliability of continuous ratings across raters. In addition, Krippendorff’s Alpha krippendorff2018content provided a robust measure suitable for ordinal scales and missing data, while the Pearson Correlation Coefficient benesty2009pearson quantified the linear consistency of expert scores. The results demonstrated substantial agreement across different measures, reinforcing the credibility of the evaluation and lending strong support to the robustness of our findings.
@@
Table 4: Comparison of explanation generation across different legal context pipelines